# Compressing Language Models for Specialized Domains

### v1

@@ Abstract
Compression techniques such as pruning and quantization offer a solution for more efficient deployment of language models (LMs), albeit with small performance drops in benchmark performance. However, general-purpose LM compression methods can negatively affect performance in specialized domains (e.g. biomedical or legal). Recent work has sought to address this, yet requires computationally expensive full-parameter fine-tuning. To this end, we propose cross-calibration, a novel training-free approach for improving the domain performance of compressed LMs. Our approach effectively leverages Hessian-based sensitivity to identify weights that are influential for both in-domain and general performance. Through extensive experimentation, we demonstrate that cross-calibration substantially outperforms existing approaches on domain-specific tasks, without compromising general performance. Notably, these gains come without additional computational overhead, displaying remarkable potential towards extracting domain-specialized compressed models from general-purpose LMs. [^1] https://github.com/mlsw/domain-compression
@@ 1 Introduction
Language models (LMs) have demonstrated remarkable performance across tasks from a range of domains (Grattafiori et al., 2024; Groeneveld et al., 2024; Yang et al., 2024). Behind this success lies a recipe with two key ingredients: highly parameterized models and extensive training. However, the vast scale of these models presents substantial challenges in their deployment and application (Treviso et al., 2023; Zhu et al., 2024). Luccioni et al. (2024) suggest that the trend towards general-purpose models has introduced substantial yet potentially unnecessary inference costs.
@@
Figure 1: Compressing large general-purpose LMs into smaller domain-specific models.
LM compression studies typically focus on preserving general-purpose performance, i.e. language modeling and commonsense reasoning capabilities (Frantar and Alistarh, 2023; Ma et al., 2023; Sun et al., 2024). However, in practice, LMs may be deployed within only one particular domain, e.g. biomedical or legal (Labrak et al., 2024; Colombo et al., 2024; Ling et al., 2024; Chen et al., 2024). This scenario unlocks new paths towards improving inference efficiency by extracting domain-specific LMs from general-purpose models (Figure 1).
Recently, Zhang et al. (2024) proposed D-Pruner, a pruning method aiming to preserve weights that are influential to both domain-specific and general capabilities. To identify such weights, D-Pruner leverages the gradient information from a composite loss function which incorporates general weight importance scores. However, this requires full-parameter fine-tuning of the LM, thus incurring substantial computational costs.
The majority of post-training LM compression methods rely upon calibration data, a small amount of data used to aid the analysis of layer activations. Recent work has shown that calibration data can impact task performance (Williams and Aletras, 2024), although Bandari et al. (2024) find that task-specific calibration data does not necessarily benefit performance. Inspired by this line of work, we investigate how to effectively leverage domain-specific data for calibration, aiming to maximize in-domain performance without sacrificing general capabilities. Our main contributions are as follows:
@@
    In contrast to previous work, we find that in-domain calibration data can play an important role in pruning, maximizing performance retention on domain-specific tasks.
@@
    We propose cross-calibration, a novel calibration method for Hessian-based compression, enabling the effective use of in-domain calibration data. Our approach is robust, with little computational overhead, yet outperforms existing general and domain-specific pruning methods.
@@ 2 Related Work
The aim of neural network pruning is to remove less important weights, therefore reducing the overall model size (LeCun et al., 1989). Pruning can be performed at the level of individual weights (unstructured), within groups of weights (semi-structured), or entire dimensions (structured) (Han et al., 2015; Mishra et al., 2021; Ma et al., 2023). In particular, 2:4 semi-structured sparsity (i.e. pruning two weights in every block of four), enables enhanced inference performance on NVIDIA GPUs (Mishra et al., 2021). However, the extensive size of LMs presents challenges in pruning them optimally (Hassibi et al., 1993). Recent work has instead decomposed LM pruning into a sequential layer-wise approach, demonstrating remarkable performance retention, even at high sparsity levels Frantar and Alistarh (2023); Sun et al. (2024); Yin et al. (2024).
@@
Early work focused on pruning deep neural networks for specific tasks (Han et al., 2015; Molchanov et al., 2017). This trend continued (Sanh et al., 2020; Lagunas et al., 2021; Kwon et al., 2022) following the advent of BERT (Devlin et al., 2019). However, the shift towards general-purpose LMs (Brown et al., 2020; Dredze et al., 2024) has led to a focus on preserving general performance (Frantar and Alistarh, 2023; Ma et al., 2023; Sun et al., 2024), i.e. language modeling and reasoning. Most recently, Zhang et al. (2024) proposed D-Pruner for domain-specific pruning, leveraging general weight importance to form a domain-specific training loss. However, this requires computationally expensive full-parameter fine-tuning, yet does not consistently outperform general-purpose pruning methods.
@@
In a post-training setting, model compression typically relies upon calibration data (Wan et al., 2024). Calibration data consists of a small number of unlabeled examples, for the generation of layer activations (Nagel et al., 2020; Hubara et al., 2021). Typically, these examples are randomly sampled from web text or pre-training datasets (e.g. C4; Raffel et al., 2020). However, recent work has illustrated the influential role that calibration data can play, impacting the downstream performance of compressed models (Williams and Aletras, 2024). Most recently, Bandari et al. (2024) suggest that pruning with calibration data based on downstream tasks does not necessarily benefit performance over generic data.
3 Cross-calibration
@@ 3.1 Preliminaries
The Optimal Brain Surgeon (OBS) (Hassibi et al., 1993) algorithm leverages second-order derivatives to accurately prune weights from a neural network. These second-order derivatives, which indicate the curvature of the loss function with respect to the weights, are organized in a square matrix known as the Hessian. Based on the Hessian $\mathbf{H}$, the OBS algorithm iteratively removes the weight $w_{m}$ with the lowest saliency $\varepsilon_{m}$, followed by applying the optimal update for the remaining weights $\boldsymbol{\delta}_{m}$:
@@
Frantar and Alistarh (2022) reformulate pruning as a case of the layer-wise compression problem, while retaining the OBS weight update procedure. Given a layer $\ell$ and input activations $\mathbf{X}_{\ell}$, the objective is to minimize the error between the original weights $\mathbf{W}_{\ell}$ and newly compressed weights $\mathbf{\widehat{W}}_{\ell}$:
$\text{argmin}_{\mathbf{\widehat{W}}}||\mathbf{X}_{\ell}\mathbf{W}_{\ell}-\mathbf{X}_{\ell}\mathbf{\widehat{W}}_{\ell}||^{2}_{2}$
As the layer input activations are derived from a fixed set of calibration data, the layer outputs (i.e. $\mathbf{Y}=\mathbf{XW}$) are also fixed. Consequently, the layer Hessian is computed as:
$\mathbf{H}=2\mathbf{X}^{\mathrm{T}}\mathbf{X}$
@@ 3.2 Motivation
A key assumption of OBS is that the original model has been sufficiently optimized, i.e. its training error is minimal. This is usually the case for LMs, which are extensively trained over large and varied corpora. Consequently, a small random sample of generic data (i.e. calibration data) proves sufficient to accurately approximate the Hessian for model compression (Frantar and Alistarh, 2023).
However, while the model weights have converged to best fit the diverse pre-training data, they also encode specialized domain-specific knowledge (Singhal et al., 2023a). We therefore question whether generic calibration data alone is sufficient to effectively isolate the weights that are crucial for domain-specific performance. As illustrated in Figure 2, certain features may be highly sensitive within a specific domain yet appear unimportant in broader, generic samples. Conversely, generic data can play a beneficial role in maintaining general capabilities like reasoning (Bandari et al., 2024).
We hypothesize that extracting domain-specific LMs while retaining general performance requires identifying weights that are influential in both cases. To this end, we propose cross-calibration (Algorithm 1), a method for Hessian-based compression that maximizes both general and in-domain performance. Our approach is twofold: (1) identifying features that are influential for specific domains, and (2) regularizing these with respect to the model, which is not converged for domain-specific data.
3.3 Mixture of Hessians
We first decompose the Hessian $\mathbf{H}$ into two distinct components: one for domain-specific features, $\mathbf{H}_{d}$, and the other, $\mathbf{H}_{g}$, for general (i.e. domain-agnostic) features. We then propose the introduction of a domain-specific calibration set $\mathcal{D}_{d}$ to approximate $\mathbf{H}_{d}$. In contrast to Zhang et al. (2024), we use in-domain data directly for assessing weight sensitivity, rather than as training data for a separate fine-tuning phase. Finally, we use a conventional generic calibration set $\mathcal{D}_{g}$ to approximate $\mathbf{H}_{g}$.
@@ 3.3 Mixture of Hessians
However, this decomposition presents two key challenges: (1) the matrix inversion required for weight updates via OBS (i.e. $\textbf{H}^{-1}$) is computationally expensive, and (2) retaining multiple Hessians substantially increases the memory requirements of compression. To address the first issue, we merge both Hessians such that matrix inversion need only be performed once. Specifically, we introduce a regularization hyperparameter $\alpha\in[0,1]$ to modulate feature influence:
$\mathbf{H}=\alpha\mathbf{H}_{d}+(1-\alpha)\mathbf{H}_{g}$
To address the second issue, we reformulate our approach as an iterative procedure, following West (1979). This enables the numerically stable computation of the Hessian while retaining only one matrix. Concretely, we consider each calibration example as a pair $(\mathbf{X}_{n},\alpha_{n})$, consisting of the input example and the regularization hyperparameter associated with the calibration set it is sampled from. This allows us to compute the Hessian iteratively as follows, where $\text{A}_{n}=\sum^{n}_{i=1}\alpha_{i}$ and $n=|\mathcal{D}_{g}\cup\mathcal{D}_{d}|$:
$\mathbf{H}_{n}=\mathbf{H}_{n-1}+\frac{\alpha_{n}}{\text{A}_{n}}(2\mathbf{X}_{n}^{\mathrm{T}}\mathbf{X}_{n}-\mathbf{H}_{n-1})$
To validate our approach, we integrate cross-calibration (CC) with SparseGPT (Frantar and Alistarh, 2023), a popular Hessian-based pruning algorithm. However, we note that cross-calibration is not tied to any specific compression algorithm (see Figure 7 in Appendix A.1 for GPTQ-M results).
@@ 4.1 Compression Methods
$\mathbf{S}_{ij}=\left[|\mathbf{W}|^{2}/\text{diag}\left((\mathbf{X}^{\mathrm{T}}\mathbf{X}+\lambda\mathbf{I})^{-1}\right)\right]_{ij}$
@@
$\mathbf{S}_{ij}=\|\mathbf{W}_{ij}|\cdot||\mathbf{X}||_{2}$
@@
D-Pruner is a domain-specific pruning method. The first step of D-Pruner is to compute the general importance $\mathbf{G}_{ij}$ using a general dataset $\mathcal{D}_{g}$, similar to SparseGPT. Secondly, it uses a composite loss function $\mathcal{L}$ to identify weights that are important for both general and domain-specific knowledge. This consists of the cross-entropy loss $\mathcal{L}_{\text{CE}}$ with a regularization term controlled by hyperparameter $\lambda_{g}$, where $\mathbf{W}^{\prime}$ is the updated weight matrix:
$\mathcal{L}\approx\mathcal{L}_{\text{CE}}+\lambda_{g}\sum_{i,j}\mathbf{G}_{ij}(\mathbf{W}^{\prime}_{ij}-\mathbf{W}_{ij})^{2}$
This is computed via stochastic gradient descent (i.e fine-tuning following Lv et al. 2024). The final saliency score is computed as follows, where $\mathcal{D}_{d}$ is a domain-specific dataset:
$\mathbf{S}_{ij}\approx\left|\frac{\partial\mathcal{L}(\mathcal{D}_{d})}{\partial\mathbf{W}_{ij}}\mathbf{W}_{ij}+\frac{1}{2}\left[\frac{\partial\mathcal{L}(\mathcal{D}_{d})}{\partial\mathbf{W}_{ij}}\mathbf{W}_{ij}\right]^{2}\right|$
GPTQ-M (Frantar et al., 2024).
@@
Algorithm 1 Cross-calibration (CC) - simplified.
Generic dataset $\mathcal{D}_{g}$, domain dataset $\mathcal{D}_{d}$, regularization $\alpha$, embedding matrix $\mathbf{U}$, and weight matrices $\mathcal{W}$.
$\mathcal{D}\leftarrow\{(x,\alpha)\mid x\in\mathcal{D}_{d}\}\cup\{(x,1-\alpha)\mid x\in\mathcal{D}_{g}\}$
for $i\leftarrow 1$ to $|\mathcal{D}|$ do
     $(x_{i},\alpha_{i})\leftarrow\mathcal{D}_{i}$
     $\mathbf{X}_{i}\leftarrow x_{i}\mathbf{U}$
end for
for $\mathbf{W}\in\mathcal{W}$ do
     $\mathbf{H}\leftarrow\mathbf{0}$, $A\leftarrow 0$
     for $i\leftarrow 1$ to $|\mathcal{D}|$ do
         $A\leftarrow A+\alpha_{i}$
         $\mathbf{H}\leftarrow\mathbf{H}+\frac{\alpha_{i}}{A}(2\mathbf{X}_{i}^{\mathrm{T}}\mathbf{X}_{i}-\mathbf{H})$
     end for
     $\mathbf{W}\leftarrow\text{Compress}(\mathbf{W},\mathbf{H})$triangleright E.g. via SparseGPT.
     for $i\leftarrow 1$ to $|\mathcal{D}|$ do
         $\mathbf{X}_{i}\leftarrow\mathbf{X}_{i}\mathbf{W}$
     end for
end for
@@
Guided by prior work (Sun et al., 2024; Frantar et al., 2024), we focus our experiments on the following settings:
@@
  • 4-bit quantization with 2:4 sparsity. Finally, we combine 2:4 sparsity with 4-bit quantization of the remaining weights, enabling up to 5.3$\times$ GPU inference speedups (Frantar et al., 2024).

### v2

@@ Abstract
Language models (LMs) excel at tasks across diverse domains, yet require substantial computational resources during inference. Compression techniques such as pruning and quantization offer a practical path towards efficient LM deployment, exemplified by their ability to preserve performance on general-purpose benchmarks. However, general-purpose LM compression methods can negatively affect performance in specialized domains (e.g. biomedical or legal). Recent work has sought to address this issue, but requires a computationally expensive full-parameter fine-tuning pipeline. To this end, we propose MixCal, a novel calibration method designed to improve the in-domain performance of compressed LMs in a post-training setting. Through extensive experimentation, we demonstrate that MixCal substantially outperforms existing approaches on domain-specific tasks and preserves general performance. Notably, these performance gains are achieved while also reducing the computational cost of LM compression. [^1] https://github.com/mlsw/domain-compression
@@ 1 Introduction
Language models (LMs) have demonstrated remarkable performance across tasks from a range of domains (Walsh et al., 2025; Guo et al., 2025; Kamath et al., 2025). Behind this success lies a recipe with two key ingredients: highly parameterized models and extensive training. However, the vast scale of these models presents substantial challenges in their deployment and application (Treviso et al., 2023; Zhu et al., 2024). Luccioni et al. (2024) suggest that the trend towards general-purpose models has introduced substantial yet potentially unnecessary inference costs.
@@
Figure 1: MixCal (§3) is applied within the calibration phase of model compression, leveraging a combination of activations from generic and domain-specific data.
LM compression studies typically focus on preserving general-purpose performance, i.e. language modeling and commonsense reasoning capabilities (Frantar and Alistarh, 2023; Ma et al., 2023; Sun et al., 2024). However, in practice, LMs may be deployed within only one particular domain, e.g. biomedical or legal (Labrak et al., 2024; Colombo et al., 2024; Ling et al., 2024; Chen et al., 2024). This scenario unlocks new paths towards improving inference efficiency by extracting domain-specific LMs from general-purpose models.
Zhang et al. (2024) proposed D-Pruner, a pruning method aiming to preserve weights that are influential to both domain-specific and general capabilities. To identify such weights, D-Pruner leverages the gradient information from a composite loss function that incorporates general weight importance scores. However, this requires a full-parameter fine-tuning pipeline for the LM, thus incurring substantial computational costs.
The majority of post-training LM compression methods rely upon calibration data, a small amount of data used to aid the analysis of layer activations. Recent work has shown that calibration data can impact task performance (Williams and Aletras, 2024; Bandari et al., 2024; Williams et al., 2025). Inspired by this line of work, we investigate how to effectively leverage domain-specific data for calibration, aiming to maximize in-domain performance without sacrificing general capabilities. Our main contributions are as follows:
@@
    We find that in-domain calibration data can play an important role in LM pruning, maximizing performance retention on domain-specific tasks.
@@
    We propose MixCal, a novel calibration method for Hessian-based compression, enabling the effective use of in-domain calibration data (Figure 1). Our approach outperforms existing general and domain-specific pruning methods while offering greater computational efficiency.
@@ 2 Related Work
The aim of neural network pruning is to remove less important weights, therefore reducing the overall model size (LeCun et al., 1989). Pruning can be performed at the level of individual weights (unstructured), within groups of weights (semi-structured), or entire dimensions (structured) (Han et al., 2015; Mishra et al., 2021; Ma et al., 2023). In particular, 2:4 semi-structured sparsity (i.e. pruning two weights in every block of four) enables enhanced inference performance on NVIDIA GPUs (Mishra et al., 2021). However, the extensive size of Transformer-based (Vaswani et al., 2017) LMs presents challenges in pruning them optimally (Hassibi et al., 1993). Recent work has instead decomposed LM pruning into a sequential layer-wise approach, demonstrating remarkable performance retention, even at high sparsity levels (Frantar and Alistarh, 2023; Sun et al., 2024; Yin et al., 2024).
@@
Early work focused on pruning deep neural networks for specific tasks (Han et al., 2015; Molchanov et al., 2017). This trend continued (Sanh et al., 2020; Lagunas et al., 2021; Kwon et al., 2022) following the advent of BERT (Devlin et al., 2019). However, the shift towards general-purpose LMs (Brown et al., 2020; Dredze et al., 2024) has led to a focus on preserving general performance (Frantar and Alistarh, 2023; Ma et al., 2023; Sun et al., 2024), i.e. language modeling and reasoning. Most recently, Zhang et al. (2024) proposed D-Pruner for domain-specific pruning, leveraging general weight importance to form a domain-specific training loss. However, this requires an expensive full-parameter fine-tuning pipeline, yet does not consistently outperform general-purpose pruning methods.
@@
In a post-training setting, model compression usually relies upon calibration data (Wan et al., 2024). Calibration data consists of a small number of unlabeled examples for the generation of layer activations (Nagel et al., 2020; Hubara et al., 2021). Typically, these examples are randomly sampled from web text or pre-training datasets (e.g. C4; Raffel et al., 2020). Recent work has illustrated the influential role that calibration data can play, impacting the downstream performance of compressed models (Williams and Aletras, 2024; Williams et al., 2025). However, Bandari et al. (2024) suggest that pruning using downstream task data does not necessarily outperform generic data on the corresponding task.
3 MixCal
@@ 3.1 Preliminaries
The Optimal Brain Surgeon (OBS; Hassibi et al., 1993) algorithm leverages second-order derivatives to accurately prune weights from a neural network. These second-order derivatives, which indicate the curvature of the loss function with respect to the weights, are organized in a square matrix known as the Hessian $\mathbf{H}$. Based on the Hessian, the OBS algorithm iteratively removes the weight $w_{m}$ with the lowest saliency $\varepsilon_{m}$, followed by applying the optimal update for the remaining weights $\boldsymbol{\delta}_{m}$:
@@
Frantar and Alistarh (2022) reformulate pruning as a case of the layer-wise compression problem, while retaining the OBS weight update procedure. Given a layer with input activations $\mathbf{X}$, the aim is to minimize the error between the original layer weights $\mathbf{W}$ and newly compressed weights $\mathbf{\widehat{W}}$:
$\underset{\mathbf{\widehat{W}}}{\arg\min}\;\|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{\widehat{W}}\|^{2}_{F}$
As the layer input activations are derived from a fixed set of calibration data, the layer outputs (i.e. $\mathbf{Y}=\mathbf{XW}$) are also fixed. Consequently, the layer-wise Hessian is computed as:
$\mathbf{H}=2\mathbf{X}^{\top}\mathbf{X}$
@@ 3.2 Motivation
Language models are trained over diverse and expansive corpora. Accordingly, a random sample of generic calibration data often proves sufficient to approximate the Hessian (Frantar and Alistarh, 2023). These generic samples can also help preserve general capabilities (Bandari et al., 2024). However, LM weights additionally encode domain-specific knowledge (Singhal et al., 2023). For example, Figure 2 illustrates that a feature may be sensitive in a target domain yet appear unimportant under generic data. Therefore, we hypothesize that compressing LMs for specialized domains while maintaining general performance requires preserving weights that are important under both settings. To this end, we propose MixCal (Algorithm 1).
3.3 Mixing Activations
First, we extend layer-wise compression to a multi-objective setting. The standard layer-wise reconstruction loss for a layer with weights $\mathbf{W}$, compressed weights $\widehat{\mathbf{W}}$, and input activations $\mathbf{X}$ is:
$\mathcal{L}(\widehat{\mathbf{W}},\mathbf{X})=\|\mathbf{X}\mathbf{W}-\mathbf{X}\widehat{\mathbf{W}}\|^{2}_{F}$
We introduce two distinct sources of calibration data, a domain-specific dataset $\mathcal{D}_{d}$, and a general-purpose dataset $\mathcal{D}_{g}$. Samples from these datasets are used to form the domain-specific and general-purpose activations, $\mathbf{X}_{d}$ and $\mathbf{X}_{g}$, respectively. We then define a combined loss, which balances reconstruction between both calibration datasets:
$\mathcal{L}_{\text{weighted}}=\alpha\mathcal{L}(\widehat{\mathbf{W}},\mathbf{X}_{d})+\beta\mathcal{L}(\widehat{\mathbf{W}},\mathbf{X}_{g})$
To control the relative contribution from each loss term, we introduce the weight coefficients $\alpha$ and $\beta$. We impose $\alpha+\beta=1$ to avoid an arbitrary scaling, setting $\beta=1-\alpha$. The loss is therefore parameterized by a single coefficient, $\alpha\in[0,1]$. This loss function yields the following Hessian:
$\mathbf{H}_{\text{weighted}}=2(\alpha\mathbf{X}_{d}^{\top}\mathbf{X}_{d}+\beta\mathbf{X}_{g}^{\top}\mathbf{X}_{g})$
However, this formulation treats the domain-specific and general-purpose activations independently. The Hessian $\mathbf{H}_{\text{weighted}}$ is a weighted sum of the domain-specific and general-purpose Hessians, namely $2\mathbf{X}_{d}^{\top}\mathbf{X}_{d}$ and $2\mathbf{X}_{g}^{\top}\mathbf{X}_{g}$. These terms only capture the second-order structure for their corresponding dataset. Consequently, $\mathbf{H}_{\text{weighted}}$ does not encode interactions between the two datasets.
@@ 3.3 Mixing Activations
We therefore propose a strategy which also incorporates such cross-dataset interactions. Concretely, we construct a mixed activation matrix $\mathbf{Z}$ that combines the activations from each of the calibration datasets. This is used to form the loss $\mathcal{L}_{\text{mix}}$:
$\mathbf{Z}=\sqrt{\alpha}\mathbf{X}_{d}+\sqrt{\beta}\mathbf{X}_{g},\quad\mathcal{L}_{\text{mix}}=\mathcal{L}(\widehat{\mathbf{W}},\mathbf{Z})$
Expanding the Hessian for $\mathcal{L}_{\text{mix}}$ illustrates the additional cross term that is introduced:
$\mathbf{H}_{\text{mix}}$ $=\mathbf{H}_{\text{weighted}}+2\sqrt{\alpha\beta}(\mathbf{X}_{d}^{\top}\mathbf{X}_{g}+\mathbf{X}_{g}^{\top}\mathbf{X}_{d})$
The contribution from this term can be interpreted as how strongly the domain-specific and general-purpose activations align.
Finally, we note a connection to the training-time data augmentation concept of mixup (Zhang et al., 2018), which forms linear combinations of pairs of examples and their labels. Mixup has previously been applied directly to activations (Verma et al., 2019) and Transformer LMs (Sun et al., 2020). While our approach also forms linear combinations of activations, we use unlabeled examples solely to approximate the Hessian for compression.
3.4 Implementation
In practice, we compute the Hessian using the form $\mathbf{H}_{\text{mix}}=2\mathbf{Z}^{\top}\mathbf{Z}$. When the total number of calibration examples is held constant, this can reduce the cost of calibration, since each update to the Hessian incorporates two examples simultaneously. We analyze the resulting efficiency gains in §5.3.
To empirically validate our approach, we integrate MixCal with SparseGPT (Frantar and Alistarh, 2023), a popular Hessian-based pruning algorithm. However, we emphasize that MixCal is not tied to any specific compression algorithm (see Figure 7, Appendix A.1 for results with GPTQ-M).
@@ 4.1 Compression Methods
$\mathbf{S}_{ij}=\left[|\mathbf{W}|^{2}/\text{diag}\left((\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\right)\right]_{ij}$
@@
$\mathbf{S}_{ij}=|\mathbf{W}_{ij}|\cdot\|\mathbf{X}\|_{2}$
@@
D-Pruner is a domain-specific pruning method. The first step of D-Pruner is to compute the general importance $\mathbf{G}_{ij}$ using a general dataset $\mathcal{D}_{g}$, similar to SparseGPT. Second, it uses a composite loss function $\mathcal{L}_{\text{DP}}$ to identify weights that are important for both general and domain-specific knowledge. This consists of the cross-entropy loss $\mathcal{L}_{\text{CE}}$ with a regularization term controlled by hyperparameter $\lambda_{g}$, where $\mathbf{W}^{\prime}$ is the updated weight matrix:
$\mathcal{L}_{\text{DP}}\approx\mathcal{L}_{\text{CE}}+\lambda_{g}\sum_{i,j}\mathbf{G}_{ij}(\mathbf{W}^{\prime}_{ij}-\mathbf{W}_{ij})^{2}$
The gradients are computed using a full-parameter fine-tuning pipeline (Lv et al., 2024). Given a domain-specific dataset $\mathcal{D}_{d}$, the saliency is:
$\mathbf{S}_{ij}\approx\left|\frac{\partial\mathcal{L}_{\text{DP}}(\mathcal{D}_{d})}{\partial\mathbf{W}_{ij}}\mathbf{W}_{ij}+\frac{1}{2}\left[\frac{\partial\mathcal{L}_{\text{DP}}(\mathcal{D}_{d})}{\partial\mathbf{W}_{ij}}\mathbf{W}_{ij}\right]^{2}\right|$
Algorithm 1 MixCal (simplified).
1:Domain dataset $\mathcal{D}_{d}$, generic dataset $\mathcal{D}_{g}$, weight $\alpha\in[0,1]$, embedding matrix $\mathbf{U}$, and weight matrices $\mathcal{W}$.
2:for $i\leftarrow 1$ to $|\mathcal{D}_{d}|$ do $\triangleright$ Since $|\mathcal{D}_{d}|=|\mathcal{D}_{g}|$.
3:  $(\mathbf{X}_{d,i},\,\mathbf{X}_{g,i})\leftarrow(\mathcal{D}_{d,i}\cdot\mathbf{U},\,\mathcal{D}_{g,i}\cdot\mathbf{U})$
4:end for
5:for $\mathbf{W}\in\mathcal{W}$ do
6:  $\mathbf{H}\leftarrow\mathbf{0}$
7:  for $i\leftarrow 1$ to $|\mathcal{D}_{d}|$ do
8:   $\mathbf{Z}_{i}\leftarrow\sqrt{\alpha}\mathbf{X}_{d,i}+\sqrt{1-\alpha}\mathbf{X}_{g,i}$
9:   $\mathbf{H}\leftarrow\frac{i-1}{i}\mathbf{H}+\frac{2}{i}(\mathbf{Z}_{i}^{\top}\mathbf{Z}_{i})$
10:  end for
11:  $\mathbf{W}\leftarrow\text{Compress}(\mathbf{W},\,\mathbf{H})$ $\triangleright$ E.g. via SparseGPT.
12:  for $i\leftarrow 1$ to $|\mathcal{D}_{d}|$ do
13:   $(\mathbf{X}_{d,i},\,\mathbf{X}_{g,i})\leftarrow(\mathbf{X}_{d,i}\mathbf{W},\,\mathbf{X}_{g,i}\mathbf{W})$
14:  end for
15:end for
GPTQ-M (Frantar et al., 2025).
@@
Guided by prior work (Sun et al., 2024; Frantar et al., 2025), we focus our experiments on the following settings:
@@
[…]